Find three consecutive even integers such that the first integer and three times the last integer is 20 less than five times the second integer.
Answer by CubeyThePenguin(3113)

consecutive even integers: (x-2), x, (x+2)

(x-2) + 3(x+2) = 5x - 20
x - 2 + 3x + 6 = 5x - 20
4x + 4 = 5x - 20
24 = x

The integers are 22, 24, and 26.